A compound consist of 65.45% C, 5.492% H, and 29.06%O on a mass basis and has a molar mass of 110 g/mol. Determine the molecular formula of the compound
1
Expert's answer
2012-09-04T10:30:35-0400
let's assume we have 1 mole of this compound. If we have 110 g, so:
mass C = 0.6545 x 110 = 72 g
mass H = .05492 x 110 = 6.05 g
mass O = .2906 x 110 = 32 g
Calculate the number of moles:
moles C = 72g / (12g/mole) = 6 moles
moles H = 6.05 / (1.01g/mole) = 6 moles
moles O = 32 g / (16 g/mole) = 2 moles

so molecular formula& is C6H6O2